OpenStudy (anonymous):

how do you differentiate ln (x+3/x-2)?

OpenStudy (anonymous):

using the quotient rule: \[\frac{ d }{ dx } \frac{ f(x) }{ g(x) } = \frac{ f'(x)g(x) - f(x)g'(x) }{ (g'(x))^2 }\]

OpenStudy (anonymous):

fprime(x)/f(x) where f(x)=(x+3)/(x-2)

OpenStudy (anonymous):

Separate into two separate logarithms. Then Integrate.